The real risk-free rate of interest is 3.8%. Inflation is expected to be 2.5% this year and 4.0% during the next 2 years. . Assume that the maturity risk premium is 0.5% for 2-year and 0.8% for 3-year period. What is the yield on 2-year Treasury securities? What is the yield on 3-year Treasury securities
1
Expert's answer
2020-06-09T15:55:40-0400



"r=r*+IP+DRP+LP+MRP"

r = required return on a security ???

r* = real risk-free rate of interest =3.8%

IP = inflation premium = 2.5% & 4.0% for next 2 yrs

DRP = default risk premium = 0 as Tresurry Security

LP = liquidity premium = 0 as Tresurry Security

MRP = maturity risk premium =0.5 for 2yrs and 0.8 for 3years


For 2 Yr yield, we have IP = "{(2.5+4)\over2}=3.25\%"

So we have"r=4\%+3.5\%+0.5\%+0+0=8\%"


For 3 Yr yield, we have IP = "{(2.5\%+4\%+4\%)\over3}=3.5\%"

So we have "r=4\%+3.5\%+0.8\%+0+0=8.3\%" ...............